Solve (x^2-3)^2-11(x^2-3)+30=0
Answer by DrBeeee(684)

Let y = x^2 -3, your quadratic reduces to
y^2 - 11y + 30 = 0 which factors into
(y-6)(y-5)=0 yielding
y = 5,6 or
x^2 - 3 = 5,6 or
x^2 = 8,9
x= 2sqrt(2),3